The root of the equation 4/5 (x+5/6)-2/3 (x-1/4)=10/9

Grade:8

1 Answers

Vikas TU
4 years ago
Hiii 
root of the equation 4/5 (x+5/6)-2/3 (x-1/4)=10/9
4/5 x + 4/5 * 5/6 - 2/3 x + 2/3 * 1/4 = 10/9 
4/5 x + 2/3 -2/3 x + 1/6 = 10/9 
12x - 10x / 15 + 1/3 ( 2+ 1/2 -10/3) =0
solve this , you will get 
2x - 25 =0 
x = 25/2 
Hope this helps